DENSITY FUNCTIONAL THEORY CALCULATION OF EXCITED-STATE PROPERTIES OF ACTINIDES

\underline{ZHIYONG ZHANG}, ROBERT J. HARRISON, P.O.Box 999, MS K1-83, Richland, WA99352.

Ground- and excited-state geometres and vibrational frequencies of UO_{2}Cl_{4}^{2−} are obtained by DFT calculations. Various approximate functionals are tested and the results are compared with experimental values. The effect of the spin-orbit interaction on the vibrational frequencies is also studied.